Nina has a total of $ 2500 to invest. She puts part of it in an investment that pays 6% per year. The remaining money she puts in an investment that pays 5% per year. At the end of one year she has earned $138.50 in interest. How much money did she put in each of the investments?
Answer by checkley79(3341)

.06x+.05(2,500-x)=138.50
.06x+125-.05x=138.50
.01x=138.50-125
.01x=13.50
x=13.50/.01
x=$1,350 aamount invested @ 6%
2,500-1,350=$1,150 amount invested @ 5%.
Proof:
.06*1,350+.05*1,150=138.50
81+57.50=138.50
138.50=138.50